The genetic map between four loci is given: P—2cM—Q-1cM-R---6cM---S

The probability of recombination between P and S is

Select one:

a. 0.009

b. 0.012

c. 0.00012

d. 0.09

e. 0.9
Answer #1

Distance between P and S is 2+1+6= 9 cM.

Probability of producing recombinants= No. of recombinants/ No. of offsprings.

Also,

Genetic map distance= No. Of recombinants/Total no. of offsprings X 100

therefore from above 2 equation we can conclude that,

Genetic map distance= probability of producing recombinantsX 100

probability of producing recombinants= 9 / 100= 0.09

Answer- d. 0.09